A concrete retaining wall of 10m long, 5m high, and 200mm thick is to
be painted with latex paint.
1. Determine the lateral area to be painted.
2. How much is the total cost if the painting cost is Php 400 per
sqm.
1
Expert's answer
2020-10-12T11:55:13-0400



Solution.

  1. The concrete retaining wall is a rectangular parallelepiped. The lateral surface of a rectangular parallelepiped consists of four rectangles: AKND, LBCM, KLBA, NMCD. The rectangles AKND and LBCM, KLBA and NMCD are equal. Let's define the areas of these rectangles:

10 * 5 = 50 m2

5 * 0,2 = 1 m2

Let's determine the area of the lateral surface of the parallelepiped, it is equal to the sum of the rectangles:

S = 50 + 50 + 1+ 1 = 102 m2

2. Let's determine the total cost:

102 * 400 = 40800 php

Answer: 1. 102 м2

2. 40800 php